noobgamer:

The perimeter of Dana's rectangular garden is 61 feet. The length of her garden is 5 feet more than twice the width. The perimeter of a rectangle can be found by adding the lengths of its four sides.

mhanifa:

Do you know perimeter formula?

Extrinix:

Ok so to start off, let's plug in all of the numbers: \(61=(2(w)+5)+(2(w)+5)+w+w\) We need to simplify it: \(61=6w+10\) Then solve for the width (w): \(61=6w+10\) subtract 10 from both sides \(51=6w\) divide by 6 on both sides \(8.5=w\) Now we need to plug it in: \(61=(2(8.5)+5)+(2(8.5)+5)+8.5+8.5\) Which we can simplify it into: \(P=2l+2w\) \(61=2(22)+2(8.5)\) So, from this we can get our length and width: \(l = 22\) \(w = 8.5\)

noobgamer:

thank you!

mhanifa:

P = 2(w + l) l = 2w + 5 61 = 2(w + 2w + 5) Solve for w: 61 = 2(3w + 5) 3w + 5 = 30.5 3w = 25.5 w = 8.5 feet Find l: l = 2w + 5 l = 2(8.5) + 5 l = 22 feet

mhanifa:

@extrinix wrote:
Ok so to start off, let's plug in all of the numbers: \(61=(2(w)+5)+(2(w)+5)+w+w\) We need to simplify it: \(61=6w+10\) Then solve for the width (w): \(61=6w+10\) subtract 10 from both sides \(51=6w\) divide by 6 on both sides \(8.5=w\) Now we need to plug it in: \(61=(2(8.5)+5)+(2(8.5)+5)+8.5+8.5\) Which we can simplify it into: \(P=2l+2w\) \(61=2(22)+2(8.5)\) So, from this we can get our length and width: \(l = 22\) \(w = 8.5\)
Sorry, didn't see your answer

Extrinix:

It's alright.